OpenStudy (usukidoll):

find some constants $a$ and $b$ (not both zero), such that $u(x,y)=f(ax+by)$ is a generic solution, where $f$ is an arbitrary $C^1$ function.\\ 8a. $u_x+2u_y=0$\\

OpenStudy (usukidoll):

\[u_x+2u_y=0 \]

OpenStudy (usukidoll):

FInd some constants\[a\] and \[b\] not both zero such that \[u(x,y) = f(ax+by)\] is a generic solution that satisfies the equation which is posted above. I want to know what the heck are they talking about?

hartnn (hartnn):

lol, even i am wondering what generic solution is and how to find it...

OpenStudy (usukidoll):

wait I'll give you the link to the book

OpenStudy (usukidoll):

I tried using example 5.... got me in a mess

OpenStudy (usukidoll):

@ganeshie8

OpenStudy (usukidoll):

any idea on how this works @hartnn?

hartnn (hartnn):

entirely new concept... trying to find and use \(u_x ,u_y\)

hartnn (hartnn):

af' +2bf' =0 a+2b =0

OpenStudy (usukidoll):

the back of the book (the one I have a hardcopy) mentioned about a ratio a/b a+2b = 0 like when a = 2 and b = -1 or something like that?!

hartnn (hartnn):

right, thats what i got just find u_x and u_y

hartnn (hartnn):

u = f(ax+by) \(u_x = f'(ax+by) (ax+by)' = af'(ax+by)\) similarly find u_y

OpenStudy (usukidoll):

wait how... and then what happens afterwards?

OpenStudy (usukidoll):

 like that?